# Upper limit for superconducting transition temperature in Eliashberg – McMillan theory

### v1

@@ Abstract
where $E_{F}=p^{2}/2m$ is Fermi energy, $v_{F}=p_{F}/m$ is electron velocity at the Fermi surface. The value of $\frac{e^{2}}{\hbar v_{F}}$, as is well known, represents the dimensionless coupling for Couloumb interaction and for typical metals it is of the order of or greater than unity. The factor of $\sqrt{\frac{m}{M}}$ determines isotopic effect.
@@
$T_{c}^{\star}\sim\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}\alpha^{2}mc^{2}=\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}\frac{me^{4}}{\hbar^{2}}=\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}Ry$ (12)
where $Ry=me^{4}/\hbar^{2}\approx$ 13.6 eV is the Rydberg constant. Here we have obtained the same combination of fundamental (atomic) constants, which was suggested in Ref. Trach , by some quite different reasoning, as determining the upper limit of superconducting critical temperature. However, our expression contains an extra factor of $r_{s}^{-1}$, which necessarily reflects the specifics of a material under consideration (density of conduction electrons), so that the value of $T_{c}^{\star}$ is in no sense universal.

### v2

@@ Abstract
where $E_{F}=p_{F}^{2}/2m$ is Fermi energy, $v_{F}=p_{F}/m$ is electron velocity at the Fermi surface. The value of $\frac{e^{2}}{\hbar v_{F}}$, as is well known, represents the dimensionless coupling for Couloumb interaction and for typical metals it is of the order of or greater than unity. The factor of $\sqrt{\frac{m}{M}}$ determines isotopic effect.
@@
$T_{c}^{\star}\sim\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}\alpha^{2}mc^{2}/2\sim\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}\frac{me^{4}}{2\hbar^{2}}\sim\frac{0.2}{r_{s}}\sqrt{\frac{m}{M}}Ry$ (12)
where $Ry=me^{4}/2\hbar^{2}\approx$ 13.6 eV is the Rydberg constant. Here we have obtained the same combination of fundamental (atomic) constants, which was suggested in Ref. Trach , by some quite different reasoning, as determining the upper limit of superconducting critical temperature. However, our expression contains an extra factor of $r_{s}^{-1}$, which necessarily reflects the specifics of a material under consideration (density of conduction electrons), so that the value of $T_{c}^{\star}$ is in no sense universal.